PROVE that N (a) is a subgroup of G Where N (a) is a normalizer of a in G


1
Expert's answer
2022-02-24T05:40:54-0500

Let us prove that a normalizer "N (a)=\{g\in G: ga=ag\}" of "a" is a subgroup of "G."

Let "h,k\in N(a)." Then "ha=ah" and "ka=ak." It follows that

"(hk)a=h(ka)=h(ak)=(ha)k=(ah)k=a(hk),"

and hence "hk\in N(a)."

It also follows from "ha=ah" that "a^{-1}haa^{-1}=a^{-1}aha^{-1}," and hence "a^{-1}h=ha^{-1}." We conclude that "a^{-1}\in N(a)."

Therefore, "N (a)" is a subgroup of "G."